OpenStudy (destinym823):

A line contains the points (3, –2) and (–6, –8). Write the equation of the line using point-slope form.

OpenStudy (imstuck):

Do you know the formula to find the slope of a line given 2 points?

OpenStudy (destinym823):

y=mx+b ?

OpenStudy (imstuck):

\[m=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (imstuck):

The formula you gave above is the slope-intercept form of a line.

OpenStudy (imstuck):

First let's find the slope using the points. Fill in the formula like this to find the slope:

OpenStudy (destinym823):

i got -10/9 is that right ?

OpenStudy (imstuck):

\[m=\frac{ -8-(-2) }{ -6-3 }\]

OpenStudy (imstuck):

Actually, no, that's incorrect. -8-(-2) is the same thing as -8+2, so the numerator, or the rise of that slope is -6

OpenStudy (imstuck):

The run, or the denominator, is -9

OpenStudy (imstuck):

So the slope is \[\frac{ -6 }{ -9 }\] which can be reduced and simplified to \[\frac{ 2 }{ 3}\]

OpenStudy (destinym823):

oh i see

OpenStudy (imstuck):

So the slope is 2/3

OpenStudy (imstuck):

The formula for point-slope is as follows:

OpenStudy (imstuck):

\[(y-y _{1})=m(x-x _{1})\]

OpenStudy (imstuck):

You found the "m", the slope, now use ONE of the given points to complete the point-slope equation for a line. If you use the first point of (3, -2), your equation will look like this: \[(y-(-2))=\frac{ 2 }{ 3 }(x-3)\]

OpenStudy (imstuck):

Simplified you get: \[(y+2)=\frac{ 2 }{ 3 }(x-3)\]

OpenStudy (imstuck):

If you use the other point, you will get: \[(y-(-8))=\frac{ 2 }{ 3 }(x-(-6))\]

OpenStudy (imstuck):

Which simplifies to \[(y+8)=\frac{ 2 }{ 3 }(x+6)\]

OpenStudy (imstuck):

I'm going to revert in part of this message to all capitals just to make a point and hopefully it will help you make some sense out of this, so please don't be offended at all caps here...

OpenStudy (destinym823):

no i wont not at all thank you so much for the help i really appreciate it a lot

OpenStudy (imstuck):

What you need to understand as a students is that EVEN THOUGH THE EQUATIONS ARE DIFFERENT FROM EACH OTHER, THEY REPRESENT THE EXACT SAME LINE BECAUSE BOTH THOSE POINTS ARE ON THE SAME LINE!!! Using one or the other will be correct.

OpenStudy (destinym823):

thank you

OpenStudy (imstuck):

Any time!!!